Circular No. 6044

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994W IN NGC 4041 A. Bragaglia, Bologna Astronomical Observatory (BAO); and U. Mu-

nari and R. Barbon, Asiago Astrophysical Observatory, report: “A spectrogram (range 460–780 nm) obtained with the BAO 1.5-m telescope (+ BFOSC + CCD) on July 31.9 UT showed a flat spectrum with strong Hα and Hβ emissions having radial velocity +1150 km/s and FWHM 1250 km/s. No prominent P-Cyg profile was observed. This appears to be a type-II supernova. On July 30.9, V = 14.2 and B = 14.5.”

Standard CCD photometry by H. Mikuž, Ljubljana, Slovenia (comparison star HR 4610, V = 6.13, B − V = +1.17): July 30.93 UT, V = 14.13 ± 0.03; 31.85, V = 14.03 ± 0.03, B − V = +0.50 ± 0.04. Visual magnitude estimates: July 31.261 UT, 13.4 (C. E. Spratt, Victoria, BC); Aug. 1.255, 13.3 (Spratt); 1.92, 13.6 (T. Vanmunster, Landen, Belgium).

AG DRACONIS R. Viotti, Istituto di Astrofisica Spaziale, Frascati, reports: “In coordi-

nation with the International Ultraviolet Explorer observations of this symbiotic star in outburst (IAUC 6039), M. Mignoli and A. Comastri (BAO) have observed the low-resolution (380–780 nm) spectrum of AG Dra on July 2.88 UT at the 1.5-m Bologna telescope (+ BFOSC/grism) at Loiano. The star displayed a high-temperature spectrum with a hot continuum, very strong Balmer and He II 468-nm emissions, and a large number of He I lines. The emission equivalent widths (in nm) were: Hα, 6.9; Hβ, 2.2; He II (468.6 nm), 1.8; and He I (587.6 nm), 0.55. The high-ionization 683-nm line is also probably present. C. Bedalo and S. Ferluga (Department of Astronomy, University of Trieste) obtained a high-resolution blue spectrum of AG Dra on July 14 with the 1.5-m telescope (+ AURELIE) of the Haute Provence Observatory. The emission-line equivalent widths appeared weaker than on July 2. Hγ displayed a P-Cyg absorption at −220 km/s and wings extending to ∼ ±400 km/s. The Hδ profile is masked by the N III multiplet-1 emissions. An earlier high-resolution spectrum of AG Dra was obtained on June 23.76 by C. Rossi (Astronomical Institute, La Sapienza University, Rome) and myself with the 1.8-m Padova Astronomical Observatory telescope (+ échelle spectrograph) at Asiago. A very rich emission spectrum was already present in these earlier phases of the AG Dra outburst, including strong He II 468.6-nm emission with extended wings, and very strong Hα emission with a P-Cyg profile.”

1994 August 2 Daniel W. E. Green